2. Cost-Volume-Profit: 15 marks Winson Cempany produces one product and has the capacity to manufacture 100,000 unlts annually. The following information is available Seling price Variable manufacturing costs Fined manufacturing costs Flxed marketing and administrative costs Variable marketing and admisistrative costs $52 per unit $24 per unit $300,000 annuaily $240,000annually $8 per unit Required: (a) Cakcuiate the number of units that need to be sold annually to break even (2 marks) (2 marks) (e) In an attempt to achieve better results in the marketplace, management has been looking (b) How many units wouid need to be sold to earn a target annual profit of $240,0007 el. This would result in an increase in variable marketing and administrative costs by $4 per unit, and L Calodate the number of units required to break even it management imglemented (2 marks) at changing the reward system for marketing, distribution and sales personn would reduce fixed marketing and distribution costs by $150,000. the changes i. At what number of sales units will the existing operation and the management propesed plan generate the same profe? (2 marks) ??L would you suggest management makes the changes suggested? Explain (3 marks) (d) Disouss how knowledge of the break-even point and the margin of safety will assist (2 marks) (e) Explain the importance of the relevant range in relation to cost behaviour when (2 marks) management in its risk assessment of business activities. considering the costs of expanding business activities.
Answer #1

1. Contribution per unit = Selling price - Variable mfg costs - Variable admin costs = 52 - 24 -8 = 20 / unit

Fixed costs = 300,000 + 240,000 = 540,000

Therefore Break even units = 540,000 / 20 = 27,000 units

2. For Target profit of 240,000 additonal units which needs to be sold = 240,000 / 20 = 12,000 units.

Hence total sales req will be = 27000 + 12000 = 39000 units

C. i) Revised contribution = 52 - 24 - 12 = 16 / unit

Revised fixed costs = 300000 + 90000 = 390000

Therefore BEP = 390,000 / 16 = 24,375 units.

iii) Yes the management can make the change, because it reduces the break even point.